# Coverage errors for Student’s $t$ confidence intervals comparable to those in Hall (1988)

### v1

@@ Abstract
Table 1 of Hall, (1988) contains asymptotic coverage error formulas for some nonparametric approximate 95% confidence intervals for the mean based on $n$ IID samples. The table includes an entry for an interval based on the central limit theorem using Gaussian quantiles and the Gaussian maximum likelihood variance estimate. It is missing an entry for the very widely used Student $t$ confidence intervals. This note makes a mild numerical correction for the Gaussian entry and provides an entry for the Student $t$ intervals. For skewness $\gamma$ and kurtosis $\kappa$, the corrected Gaussian formula is $3.420.14\kappa-2.16\gamma^{2}-$ 3.42 and the formula for the $t$ intervals is $0.14\kappa-2.16\gamma^{2}$. The impetus to revisit this estimate arose from the surprisingly robust performance of Student’s t statistic in randomized quasi-Monte Carlo sampling.
@@ 1 Introduction
Table 1 of Hall, (1988) gives expressions for the coverage error in several nonparametric confidence interval methods for the mean of a real valued random variable $X$ based on an IID sample of $n$ realizations of that random variable. It is missing an entry for the usual confidence interval based on Student’s $t$ statistic. It does have an entry for the case where the $t$ threshold is replaced by one from the standard Gaussian distribution and the usual unbiased sample variance is replaced by the maximum likelihood estimate (MLE). This note uses an expression from page 949 of that paper to develop a formula for the Student $t$ intervals. There is also a minor numerical correction for the Gaussian intervals.
@@
For $n\geqslant 2$, let $X_{1},\dots,X_{n}$ be IID random variables with the same distribution as $X$. Then for $\alpha<1/2$, a standard Student’s $t$ based confidence interval for $\mu$ with nominal coverage $1-2\alpha$ takes the form
@@
Table 1 of Hall, (1988) includes asymptotic coverage error formulas for ‘Norm’ and some bootstrap confidence intervals. For the necessary assumptions on the distribution of $X$, see Hall, (1988). These table entries are for $\alpha=0.025$ corresponding to the common default of 95% confidence. The table’s expression for coverage error of (2) is
$\frac{1}{n}\bigl{(}0.14\kappa-2.12\gamma^{2}-3.35)\varphi(0.975)+O(n^{-2})$ (3)
where $\varphi$ is the $\mathcal{N}(0,1)$ probability density function, and equation (3) combines equation (4.15) and the row labeled ‘Norm’ in Table 1 of Hall, (1988).
@@
Section 2 rederives and mildly corrects Hall’s formula for coverage errors in 95% confidence intervals based on the Gaussian interval. Section 3 then constructs formulas for the Student $t$ interval. The key step is to use an asymptotic formula from Abramowitz and Stegun, (1972) relating quantiles of Student’s $t$ distribution to those of the $\mathcal{N}(0,1)$ distribution. Both tasks were made simpler by noticing that Hall’s derivation for the ‘Norm’ interval does not use all of the machinery that the bootstrap intervals in that paper require.
@@ 2 Derivation
$1-2\alpha-2n^{-1}z_{1-\alpha}\Bigl{(}\frac{1}{12}\kappa(z_{1-\alpha}^{2}-3)\\ -\frac{1}{18}\gamma^{2}\bigl{(}z_{1-\alpha}^{4}+2z_{1-\alpha}^{2}-3\bigr{)}-\frac{1}{4}\bigl{(}z_{1-\alpha}^{2}+3\bigr{)}\Bigr{)}\varphi(z_{1-\alpha})+O(n^{-3/2}).$
@@
$\frac{z_{1-\alpha}}{n}\Bigl{(}\frac{z_{\alpha}^{2}-3}{6}\kappa-\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{9}\gamma^{2}-\frac{z_{\alpha}^{2}+3}{2}\Bigr{)}\varphi(z_{\alpha})+O(n^{-3/2})$ (5)
where even powers of $z_{1-\alpha}$ and $z_{\alpha}$ are equal because $z_{\alpha}=-z_{1-\alpha}$. The corresponding entry for Table 1, which leaves out the factor $z_{1-\alpha}\varphi(z_{\alpha})/n$ is then
$\frac{z_{\alpha}^{2}-3}{6}\kappa-\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{9}\gamma^{2}-\frac{z_{\alpha}^{2}+3}{2}.$ (6)
Using $z_{1-\alpha}=z_{0.975}\doteq 1.95996$, we get $3.420.14\kappa-2.16\gamma^{2}-$ 3.42 after rounding. Hall’s Table 1 has $3.350.14\kappa-2.12\gamma^{2}-$ 3.35 which is close. He may have used $z_{1-\alpha}=1.92$ for the constant term and $z_{1-\alpha}=1.95$ for the coefficient of $\gamma^{2}$. Also, $\varphi(z_{0.975})\doteq 0.058$.
@@
$-\frac{z_{\alpha}^{2}+3}{2n}\varphi(z_{\alpha})+O(n^{-3/2}).$

### v2

@@ Abstract
Table 1 of Hall, (1988) contains asymptotic coverage error formulas for some nonparametric approximate 95% confidence intervals for the mean based on $n$ IID samples. The table includes an entry for an interval based on the central limit theorem using Gaussian quantiles and the Gaussian maximum likelihood variance estimate. It is missing an entry for the very widely used Student’s $t$ confidence intervals. This note develops such a formula. The impetus to revisit this issue arose from the surprisingly robust performance of confidence intervals based on Student’s t statistic in randomized quasi-Monte Carlo sampling. Hall’s table had $3.350.14\kappa-2.12\gamma^{2}-$ 3.35 for normal theory intervals; the corresponding entry for Student’s $t$ is $0.14\kappa-2.12\gamma^{2}$.
An earlier version of this note reported that it corrected some coverage error formulas in Hall, (1988). Two-sided errors take the form $2\Phi^{-1}(0.975)(A\kappa+B\gamma^{2}+C)\varphi(1.96)/n+O(1/n^{3/2})$ where the error may well be $O(n^{-2})$. Hall’s table showed $\Phi^{-1}(0.975)(A\kappa+B\gamma^{2}+C)$. The version intended as a correction had $2(A\kappa+B\gamma^{2}+C)$, wider by about $2/1.96\doteq 1.02$. So, Hall’s table really is proportional to the two-sided coverage errors.
@@ 1 Introduction
Table 1 of Hall, (1988) gives expressions for the coverage error in several nonparametric confidence interval methods for the mean of a real-valued random variable $X$ based on an IID sample of $n$ realizations of that random variable. It is missing an entry for the usual confidence interval based on Student’s $t$ statistic. It does have an entry for the case where the $t$ threshold is replaced by one from the standard Gaussian distribution and the usual unbiased sample variance is replaced by the maximum likelihood estimate (MLE). This note uses an expression from page 949 of that paper to develop a formula for the Student $t$ intervals.
@@
For $n\geqslant 2$, let $X_{1},\dots,X_{n}$ be IID random variables with the same distribution as $X$. Then for $20<\alpha<1/$ 2, a standard Student’s $t$ based confidence interval for $\mu$ with nominal coverage $1-2\alpha$ takes the form
@@
Table 1 of Hall, (1988) includes asymptotic coverage error formulas for ‘Norm’ and some bootstrap confidence intervals. For the necessary assumptions on the distribution of $X$, see Hall, (1988). These table entries are for $\alpha=0.025$ corresponding to the common default of 95% confidence. The coverage error of (2) for $\alpha=0.025$ is
$\frac{2}{n}\Bigl{(}0.14\kappa-2.12\gamma^{2}-3.35\Bigr{)}\varphi(1.96)+O(n^{-2}),$ (3)
where $\varphi$ is the $\mathcal{N}(0,1)$ probability density function, and the constants have been rounded. The row labeled ‘Norm’ in Table 1 of Hall, (1988) contains $3.350.14\kappa-2.12\gamma^{2}-$ 3.35, omitting the factor $2\varphi(1.96)/n$ common to all of the coverage error expressions.
@@
Section 2 rederives Hall’s formula for coverage errors in 95% confidence intervals based on the Gaussian interval. Section 3 then constructs formulas for the Student $t$ interval. The key step is to use an asymptotic formula from Abramowitz and Stegun, (1972) relating quantiles of Student’s $t$ distribution to those of the $\mathcal{N}(0,1)$ distribution. Both tasks were made simpler by noticing that Hall’s derivation for the ‘Norm’ interval does not use all of the machinery that the bootstrap intervals in that paper require.
@@ 2 Derivation
$1-2\alpha+2n^{-1}z_{1-\alpha}\Bigl{(}\frac{1}{12}\kappa(z_{1-\alpha}^{2}-3)\\ -\frac{1}{18}\gamma^{2}\bigl{(}z_{1-\alpha}^{4}+2z_{1-\alpha}^{2}-3\bigr{)}-\frac{1}{4}\bigl{(}z_{1-\alpha}^{2}+3\bigr{)}\Bigr{)}\varphi(z_{1-\alpha})+O(n^{-3/2}).$
@@
On page 949, Hall, (1988) writes: “The term $-3.35$ appearing in the third column of Table 1 arises from the difference between the standard normal distribution function and an expansion of Student’s t distribution function.” This suggests that all we must do is drop the intercept from the expression for the normal theory result, but for the avoidance of doubt we work through the expansion here to verify that this is the case.
@@
$\frac{2z_{1-\alpha}}{n}\Bigl{(}\frac{z_{\alpha}^{2}-3}{12}\kappa-\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{18}\gamma^{2}-\frac{z_{\alpha}^{2}+3}{4}\Bigr{)}\varphi(z_{\alpha})+O(n^{-3/2})$ (5)
where even powers of $z_{1-\alpha}$ and $z_{\alpha}$ are equal because $z_{\alpha}=-z_{1-\alpha}$. The corresponding entry for Table 1, which leaves out the factor $2\varphi(z_{\alpha})/n$ is then
$z_{1-\alpha}\frac{z_{\alpha}^{2}-3}{12}\kappa-z_{1-\alpha}\frac{z_{\alpha}^{4}+2z_{\alpha}^{2}-3}{18}\gamma^{2}-z_{1-\alpha}\frac{z_{\alpha}^{2}+3}{2}.$ (6)
Using $z_{1-\alpha}=z_{0.975}\doteq 1.96$ we get $3.350.14\kappa-2.12\gamma^{2}-$ 3.35 after rounding, matching Hall’s Table 1.
@@
$-\frac{2\varphi(z_{\alpha})}{n}z_{1-\alpha}\frac{z_{\alpha}^{2}+3}{2}+O(n^{-3/2})\doteq-\frac{0.78}{n}.$
@@ 3 Adjusted interval
The coefficients of $\kappa$ and $\gamma^{2}$ matches those of equation (5) for the normal theory method and the intercept term is zero.